### v3

# Motion-Zero: Zero-Shot Moving Object Control Framework for Diffusion-Based Video Generation

Abstract
Recent large-scale pre-trained diffusion models have demonstrated a powerful generative ability to produce high-quality videos from detailed text descriptions. However, exerting control over the motion of objects in videos generated by any video diffusion model is a challenging problem. In this paper, we propose a novel zero-shot moving object trajectory control framework, Motion-Zero, to enable a bounding-box-trajectories-controlled text-to-video diffusion model. To this end, an initial noise prior module is designed to provide a position-based prior to improve the stability of the appearance of the moving object and the accuracy of position. In addition, based on the attention map of the U-net, spatial constraints are directly applied to the denoising process of diffusion models, which further ensures the positional and spatial consistency of moving objects during the inference. Furthermore, temporal consistency is guaranteed with a proposed shift temporal attention mechanism. Our method can be flexibly applied to various state-of-the-art video diffusion models without any training process. Extensive experiments demonstrate our proposed method can control the motion trajectories of objects and generate high-quality videos.
Figure 1: Our Motion-Zero framework endows any pre-trained video diffusion model with the capability to manipulate object trajectories directly, circumventing the need for supplementary training. By designating the target entity in the input prompts and a sequence of bounding boxes, users can intuitively direct the motion path of the object within the generated video sequence.
1 Introduction
In recent years, the generative capabilities of diffusion models have been widely recognized in both text-to-image Dhariwal and Nichol (2021); Ho et al. (2020); Rombach et al. (2022); Song et al. (2020) and text-to-video domains Guo et al. (2023); Ho et al. (2022). Although these video models are capable of producing high-definition, high-resolution, and fluid video animations, the dynamic motion trajectories of the generated objects are relatively random in existing text-to-video models Blattmann et al. (2023); Zhang et al. (2023a). Thus, accurate control of an object’s motion trajectories in a generated video remains rudimentary.
To address this problem, various efforts have been proposed in different strategies. Some methods Zhang et al. (2023b); Wu et al. (2023a) try to control the trajectory of the moving object by providing a detailed conditional motion sequence, such as a dancing skeleton. However, the cost of acquiring conditional control sequences is non-negligible and inhibits the diverse generation of video outputs by the user. To obtain variety over the generated video with more control capability, i.e. trajectories of moving objects, several methods Wang et al. (2023c, b); Yin et al. (2023) harness a substantial dataset of motion and trajectory pairs to train the baseline models. Although the results are very impressive and promising, these methods are solely applicable to the base models on which they were trained and cannot be applied to other models directly. Furthermore, all of these methods require extensive training and significant computational resources, preventing ordinary users from using them.
Nevertheless, the fundamental text-to-video diffusion models Blattmann et al. (2023); Zhang et al. (2023a) have undergone an extremely large scale training dataset Bain et al. (2021); Xue et al. (2022). Theoretically, the pre-trained model should inherently have learned extensive knowledge about the dynamics of a variety of object movements. However, the semantics of the latent space in different modules of a video diffusion model are not explicitly defined, which makes it difficult to manually control generated video in desired motion dynamics. To some extent, methods mentioned above try to align and define those latent spaces in the baseline model by training with additional labeled datasets, and successfully harness the intrinsic knowledge of pre-trained models to control the motion. In addition, we observe that the initial noise plays a significant inspirational role in the generation of the videos. The same initial noise tends to produce videos with similar content. With these two characteristics, we can manipulate the trajectory of moving objects.
In this paper, we propose a zero-shot framework, Motion-Zero, that can be easily applied to any pre-trained video diffusion model and achieve universality and plug-and-play capabilities of controlling. To this end, we designed an initial noise prior module to provide noises based on the motion trajectories given by the user. Inspired by the fact that the values of the cross-attention map within the U-Net largely determine the generation location of the subject mentioned in the prompt Xie et al. (2023); Epstein et al. (2023), spatial constraints with an attention similarity loss over these attention maps are proposed to achieve precise manipulation of object positioning within individual frames. Furthermore, based on our observations, we find that merely imposing spatial constraints on the positions of objects can negatively impact the overall coherence and video quality. To preserve the continuity throughout the video sequence, a novel shift temporal attention mechanism is proposed to further maintain the baseline model focus on the same objects across the time axis. Through the implementation of these modules, our framework can effectively generate high-quality video controlled by bounding-box trajectories without any training process. Samples generated by Motion-Zero can be viewed in Fig.1. The main innovations of our work can be summarized as follows,
  • We propose a zero-shot framework Motion-Zero which is capable of controlling the motion trajectories of arbitrary objects within a pre-trained video generation diffusion model. Our Motion-Zero is plug-and-play and without any additional training.
  • By recognizing the influence of initial noise on video generation, we implemented an initial noise conditioning module that establishes advantageous starting conditions for controllable video diffusion.
  • To further accurate control over the generative video, we propose spatial constraints and a novel shift temporal attention mechanism aimed at ensuring the positional accuracy and temporal continuity and consistency of the target, respectively.
  • Extensive experiments demonstrate that our proposed work can be applied to any pre-trained video diffusion models. Our results can outperform those from extensively pre-trained trajectory control models.
2 Related Works
2.1 Text-To-Video Diffusion Models
As text-to-image diffusion models have flourished and demonstrated impressive results, the body of work related to text-to-video diffusion has gradually expanded. Researchers and developers are leveraging the principles of text-to-image models to extend their capabilities into the dynamic and temporally complex domain of video generation. In the realm of text-to-video (T2V) synthesis, the field has witnessed several pioneering approaches. Initial endeavors such as Imagen Video Ho et al. (2022) and Make-A-Video Singer et al. (2022) tackled T2V generation at the pixel level, leading to notable advancements but also encountering constraints in video duration and quality due to computational intensity. To address these limitations, MagicVideo Zhou et al. (2022) emerged, introducing a novel autoencoder trained specifically on video data, which, akin to the impact of Latent Diffusion Models (LDMs) Rombach et al. (2022) in the text-to-image (T2I) sector, significantly improved the computational efficiency of T2V tasks. In the current landscape, there exists an abundance of generalizable open-domain text-to-video generation models capable of producing high-definition videos with dynamic visual fidelity Blattmann et al. (2023); Zhang et al. (2023a); Mullan et al. (2023); Sterling (2023); Chen et al. (2023); Wang et al. (2023a). The models often generate either stochastic motion paths or demonstrate relatively minor variations in dynamics, underscoring the potential benefits of investigating methodologies for manually stipulating more deliberate and varied motion trajectories.
2.2 Controllable Video Generalization
There are now pre-trained methodologies Wang et al. (2023c, b); Yin et al. (2023) in place that have the capability to control the motion trajectories of objects within generated videos. VideoComposer Wang et al. (2023b) employs a two-stage training strategy to incrementally incorporate temporal information and control signals. A motion condition encoder is proposed for training, designed to enhance the model’s ability to understand and integrate motion-specific information. DragNUMA Yin et al. (2023) also employs a comparable training strategy and utilizes optical flow as a conditioning mechanism for trajectory modeling. MotionCtrl Wang et al. (2023c) synthesizes the approaches of these two works and introduces a camera control module and a trajectory control module to further refine the management of movement within generated video content. These models typically rely on extensive training on large-scale datasets such as WebVid-10M Bain et al. (2021) and HD-VILA-100M Xue et al. (2022) which lead to expensive training costs. Also, these models are typically constrained to operate on the specific models they were trained on, lacking the flexibility to interchange base video diffusion models. This limitation hinders their adaptability and limits the scope of their application to only those scenarios for which they were explicitly designed. Our model stands out in that it can be applied to any base diffusion model without the necessity for further training. This versatility not only broadens its applicability but also results in substantial savings in training costs and resources, offering a significant advantage over more rigid, model-specific methods. A contemporaneous work Ma et al. (2023) parallels our approach with a similar zero-shot setting. However, they ignore the importance of the inital noise and the control of spatial and temporal consistency.
Figure 2: Overview of our Motion-Zero. The total pipeline is shown on (a). Given the box condition $\mathcal{B}$ and the prompt condition, we generate the prior latents $\mathbf{z}_{T}$ by our Initial Noise Prior Module (INPM) as shown on (b). At timestep $t$, $\mathbf{z}_{t}$ is firstly optimized to $\mathbf{z}_{t}^{\prime}$ by the Spatial Constraints (SC). Subsequently, $\mathbf{z}_{t}^{\prime}$ is passed to the Unet with Shift Temporal Attention Module (STAM) as demonstrated on (c). All the parameters of the video diffusion are frozen. $T_{1}$ represents the number of timesteps during which SC and STAM are applied, and $T_{2}$ denotes the number of timesteps where the original video diffusion process is utilized.
3 Preliminaries
Video Diffusion Model: Video diffusion models are designed to produce high-quality and diverse videos, guided by text prompts. To save the computational costs, Rombach et al. (2022) utilize a U-Net as a denoising model within a latent space, significantly reducing the computational load in terms of both time and space.
In detail, these models employ a Variational Autoencoder (VAE), which comprises an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder compresses the original video from pixel space into a latent representation, and the decoder reconstructs the video from this latent space back to pixel space. The 3D U-Net typically consists of a series of down-sampling blocks, middle blocks, and up-sampling blocks. Each block is equipped with convolutional layers, spatial transformers, and temporal transformers. The optimization of the 3D U-Net (denoted as $\epsilon_{\theta}$) is executed through a noise prediction loss function:
$\mathcal{L}=\mathcal{E}_{\mathbf{z}_{0},c,\epsilon\sim N(0,I),t}[||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})||^{2}_{2}],$ (1)
where $z_{0}$ is the latent code of the training videos, c is the text prompt condition, $\epsilon$ is the Gaussian noise added to the latent code, and $t$ is the time step. The noised latent code $\mathbf{z}_{t}$ is determined as:
$\mathbf{z}_{t}=\sqrt{\overline{a}_{t}}\mathbf{z}_{0}+\sqrt{1-\overline{a}_{t}}\mathbf{\epsilon},\overline{a}_{t}=\prod^{t}_{i=1}a_{t},$ (2)
where $a_{t}$ is a hyper-parameter used for controlling the noise strength based on time $t$.
Motion Trajectory Control: Based on video diffusion, the task of motion trajectory control is to precisely control the motion trajectory of objects in generated videos. The optimization objective can be formulated as:
$\mathcal{L}=\mathcal{E}_{\mathbf{z}_{0},c,\epsilon\sim N(0,I),t}[||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c},\mathcal{B})||^{2}_{2}].$ (3)
Specifically, users can input a text condition $\mathbf{c}$ along with a sequence of rectangular boxes =$\mathcal{B}=\{(x_{1}^{f},y_{1}^{f}),(x_{2}^{f},y_{2}^{f})\}^{N_{f}}$, where $(x_{1}^{f},y_{1}^{f}),(x_{2}^{f},y_{2}^{f})$ are the upper left and lower right points of the box in frame $f$, $N_{f}$ is the total frame number. The boxes $\mathcal{B}$ correspond to an the position of the moving object.
4 Methodology
4.1 Overview
Existing trajectory control methods require large-scale video training data to optimize Eq.3 leading to a high computational cost. Differently, our work falls into the setting of zero-shot motion trajectory control. To this end, our Motion-Zero framework operates entirely during the inference stage, thus eliminating the need for any training and making it compatible with any pre-trained video diffusion model.
In the following section, we provide a detailed presentation of the proposed Motion-Zero framework and its individual components. The pipeline is shown in Fig.2 (a). Sec.4.2 introduces the designed initial noise prior module. Sec.4.3 is followed to describe how we constrain the moving object position and spatial consistency. Finally, Sec.4.4 details our shift temporal attention mechanism which improves the temporal consistency.
4.2 Initial Noise Prior Module
According to the theory of DDIM Inversion Ho et al. (2020), the initial noise has a significant impact on the final generated outcome. The Initial Noise Prior Module (INPM) is designed to leverage this property to provide a strong prior for the position of the moving object. Several steps are involved to integrate a moving object into a sequence of frames with a coherent prior as shown in Fig.2 (b). Firstly, a meta video $V_{meta}$ is sampled using the original Video Diffusion with spatial constraints (introduced in Sec.4.3), $\mathbf{z}^{*}\sim\mathcal{N}(0,\mathbf{I})$ as latent input, conditioned on a given prompt $\mathbf{c}$ and the first frame box $\mathcal{B}^{0}$. Then, a video latent $\mathbf{z}_{meta}$ is generated based on $V_{meta}$ from Encoder $\mathcal{E}$. Once $\mathbf{z}_{meta}$ is prepared, we perform a DDIM Inversion to obtain the corresponding noise latent representation $\mathbf{z}_{I}$. We crop the latent representation within the box $\mathcal{B}^{0}$ for each frame, creating a sequence of latent patches. Subsequently, we use a local mixup operation Zhang et al. (2017) to mix the latent patches and the initial noise $\mathbf{z}^{*}$ in the range of $\mathcal{B}^{f}$ frame by frame. This method allows us to set a coherent prior in the corresponding object’s position in the initial noises. It also ensures that the animated object maintains consistency in appearance and movement across the video frames, without incurring additional computational costs during the generation process. Details are shown in Alg.1.
Algorithm 1 Initial Noise Prior Module
Input: $\mathbf{c}$, $\mathcal{B}$
Parameter: $N_{f}$, $\lambda_{p}$
Function: $\mathbf{Pixel}(\mathbf{a},\mathbf{b})$ means the elements get from tensor $\mathbf{a}$ in the range of box $\mathbf{b}$
Output: Initial noise $\mathbf{z}^{*}$
1:  $\mathbf{z}^{*}\sim\mathcal{N}(0,\mathbf{I})$$\triangleright$ random sample the first latent code
2:  $\mathbf{V}_{meta}\leftarrow\mathbf{VideoDiffusion}(\textbf{z}^{*},\textbf{c},\mathcal{B}_{0})$$\triangleright$ Video Diffusion using SC, the first box and c as condition
3:  $\mathbf{z}_{meta}\leftarrow\mathbf{Encoder}(\mathbf{V}_{meta})$
4:  $\mathbf{z}_{I}\leftarrow\mathbf{DDIM\_Inverse}(\textbf{z}_{meta})$$\triangleright$ inverse the $\mathbf{z}_{meta}$
5:  $\textbf{z}_{T}$leftarrow $\textbf{z}^{*}$
6:  for all f=1,2,…,$N_{f}$ do
[…]
8:  end for
9:  return $\mathbf{z}_{T}$$\triangleright$ initial latent with position prior
4.3 Spatial Constraints with Attention Map
The INPM alone is insufficient for precise manipulation of an object’s trajectory. To further improve the capability of control, we introduce Spatial Constraints (SC) deployed at each denoising step $t$ to optimize the intermediate latent representation $\mathbf{z}_{t}$. This optimization is crucial for enhancing the accuracy of the moving object position and preserving spatial consistency. Within the conditional denoising architecture, cross-attention serves as the pivotal bridge that connects the text prompt with the content generated. During the denoising steps, conditioned on the prompt $\mathbf{c}$ and the intermediate features $\mathbf{z}_{t}$, the corresponding cross-attention map can be abtained as $\mathbf{A}$:
$\mathbf{A}=\mathbf{Softmax}(\mathbf{QK}^{\top}/\sqrt{d}),$ (4)
$\mathbf{Q}=\mathbf{W}_{Q}\mathbf{z}_{t},\mathbf{K}=\mathbf{W}_{K}\mathbf{c},$ (5)
where $\mathbf{Q},\mathbf{K}$ are the query and key of the $\mathbf{z}_{t}$ and $\mathbf{c}$, respectively. $\mathbf{W}_{Q},\mathbf{W}_{K}$ are two learnable matrices, which are frozen in our settings. Assuming a maximum number $N_{p}$ of prompt tokens $\{\mathbf{p}_{1},...,\mathbf{p}_{N_{p}}\}$ in condition $\mathbf{c}$, at time step $t$, it is feasible to derive $N_{p}$ cross-attention maps $\{\mathbf{A}_{1},...,\mathbf{A}_{N_{p}}\}$.
When a user specifies the index $k$ of prompt tokens $\mathbf{p}_{k}$ intended to control the trajectory of an object, along with the box $\mathcal{B}$, the box-loss can be employed to ensure that our model controls the object to appear within the input box in every frame. Specifically, the box-loss is achieved through an optimization approach that maximizes the values of the $\mathbf{A}_{k}$ for the corresponding prompt tokens inside the box, while minimizing the values outside the box. Frame by frame, we scale down the user-specified coordinate box to the corresponding coordinates in the latent space and construct a mask $\mathbf{M}^{f}$ where $1$ for areas in the box and 00 otherwise. To ensure the values inside the box for attention are maximized, we propose an intuitive solution ${\mathcal{L}_{i}}$ as
${\mathcal{L}_{i}}^{f}=1-\frac{1}{P}\sum\mathbf{g}(\mathbf{A}_{k}^{f}\cdot\mathbf{M}^{f},P),$ (6)
where $\mathbf{g}(\cdot,P)$ means the top $P$ highest value will be selected. If we employ all the values within the attention maps, it could disrupt the stability of the denoising process. Conversely, utilizing too few values could result in a less pronounced control effect. On the other hand, we aim for the attention values outside the box to be as minimal as possible which is formulated as
${\mathcal{L}_{o}}^{f}=\frac{1}{P}\sum\mathbf{g}(\mathbf{A}_{k}^{f}\cdot(1-\mathbf{M}^{f}),P).$ (7)
Within these two constraints, we can ensure that the object generated in each frame is contained within the box. However, there is no guarantee that the object will be positioned at the center of the box. To mitigate this, we propose another center-loss $\mathcal{L}_{c}$ to encourage the centroid of the $\mathbf{A}_{k}^{f}$ to closely align with the center of the box as
$[W_{\mathbf{A}_{k}^{f}},H_{\mathbf{A}_{k}^{f}}]=\frac{1}{\sum_{w,h}\mathbf{A}_{k,w,h}^{f}}[\sum_{w,h}w\cdot\mathbf{A}_{k,w,h}^{f},\sum_{w,h}h\cdot\mathbf{A}_{k,w,h}^{f}],$
${\mathcal{L}_{c}}^{f}=||(\frac{x_{1}^{f}+x_{2}^{f}}{2},\frac{y_{1}^{f}+y_{2}^{f}}{2})-(W_{\mathbf{A}_{k}^{f}},H_{\mathbf{A}_{k}^{f}})||_{1},$
where $\mathbf{A}_{k,w,h}^{f}$ is the value of $\mathbf{A}_{k}^{f}$ at the position of $(w,h)$, $(W_{\mathbf{A}_{k}^{f}},H_{\mathbf{A}_{k}^{f})}$ is the centroid position of $\mathbf{A}_{k}^{f}$. After establishing robust control over the position of the object, we recognize unexpected variations in the object’s appearance due to the substantial extent of movement. To ensure appearance consistency across frames, we strive to maintain the uniformity of the $\mathbf{A}_{k}^{f}$ within the box. To this end, we introduce a similarity loss $\mathcal{L}_{s}$ as
${\mathcal{L}_{s}}=1-\frac{1}{N_{f}-1}\sum_{f}^{N_{f}-1}\mathbf{Sim}(\mathbf{Pixel}(\mathbf{A}_{i}^{f},\mathcal{B}^{f}),\mathbf{Pixel}(\mathbf{A}_{i}^{f+1},\mathcal{B}^{f+1})),$
where $\mathbf{Sim}(\cdot,\cdot)$ means the similarity of two elements and $\mathbf{Pixel}()$ is a function getting corresponding elements of the map $\mathbf{A}^{f}$ within a box range $\mathcal{B}^{f}$. Cosine similarity is adopted here. At each timestep, the overall spatial constraints $\mathcal{L}_{sp}$ are formulated as follows:
${\mathcal{L}_{sp}}=\sum_{f}(\lambda_{i}{{\mathcal{L}_{i}}^{f}+\lambda_{o}\mathcal{L}_{o}}^{f}+\lambda_{c}{\mathcal{L}_{c}}^{f})+\lambda_{s}{\mathcal{L}_{s}},$
where $\lambda_{i},\lambda_{o},\lambda_{c},\lambda_{s}$ are hyper-parameters. By minimizing and calculating the gradient of the Eq.11, we can optimize our latent $\mathbf{z}_{t}$ in Eq.2 as follows: S
$\mathbf{z}_{t}^{\prime}\leftarrow\mathbf{z}_{t}-\beta_{t}\cdot\nabla{\mathcal{L}_{sp}},$ (12)
where $\beta_{t}$ linearly decays at each timestep $t$. Specifically, before denoising with the U-Net at each timestep $t$, we update $\mathbf{z}_{t}$ to $\mathbf{z}_{t}^{\prime}$ using Eq.12. Then, we continue the denoising process using $\mathbf{z}_{t}^{\prime}$. Under the combined effect of the aforementioned constraints, the latent variable $\mathbf{z}_{t}$ at each timestep will gradually shift towards generating high-response attention in the specified position, while ensuring that the appearance attributes of the object within the box remain unchanged. Consequently, the target object is synthesized within the bounding box area provided by the user.
4.4 Shift Temporal Attention Mechanism
After the spatial constraints on attention maps, pre-trained video diffusion models still encounter difficulties in generating sequences of continuous actions. Within the temporal module of the diffusion process, the latent representation is reshaped into the following configuration (using einops Rogozhnikov (2021) notation):
[…]
To overcome this inconsistency, we propose a Shift Temporal Attention Mechanism (STAM) to improve the dynamics of the moving object in different frames. Specifically, we shift the elements of $\mathbf{z}^{f}$ inside the $\mathcal{B}^{f}$ range with the elements inside the $\mathcal{B}^{0}$ range, as shown in Fig.2 (b). Therefore, the subsequent frames within the box range can be aligned with the box range of the first frame. The steps are shown as follows,
$\begin{split}\mathbf{z}_{w}^{f}&=\mathbf{Shift}(\mathbf{z}^{f},\mathcal{B}^{f},\mathcal{B}^{0}),\\ {\mathbf{z}_{w}}^{\prime}&=\mathbf{TemporalAttention}(\mathbf{z}_{w}),\\ {\mathbf{z}^{f}}^{\prime}&=\mathbf{Shift}({\mathbf{z}_{w}^{f}}^{\prime},\mathcal{B}^{0},\mathcal{B}^{f}),\\ \end{split}$ (13)
where $\mathbf{Shift}(\cdot,a,b)$ is the shift operation, ${\mathbf{z}^{f}_{w}}$ means the shifted latent ${\mathbf{z}^{f}}$ of frame $f$ and ${\mathbf{z}_{w}}=[{\mathbf{z}^{1}_{w}},..,{\mathbf{z}^{f}_{w}}]$.By applying STAM to a $\mathbf{TemporalAttention}$ layer, we can achieve coherence in the motion of moving objects without additional training, and without incurring extra computational costs during inference.
5 Experiments
5.1 Implementation Details
Figure 3: Quality comparison results on different methods. We take one frame from every three frames. The input prompt of left part: A fish is swimming in the sea. Right part: A bear is walking on the grass. We employed ModelScope (a) and ZeroScope (b) as our baseline models and compared the effect of incorporating additional prompts with the integration of our Motion-Zero. In addition, we conducted a comparative analysis with the pre-trained MotionCtrl model (c). The red dots serve as the control conditions. Zoom in for the best view.
Our algorithm is fully implemented during the inference stage, thus it does not require any training. The hyper-parameter $\lambda_{i},\lambda_{o}$ are set to 1, $\lambda_{c}$ is set to 0.05, $\lambda_{s}$ is set to 0.5, $\lambda_{p}$ is set to 0.8. To balance the trade-off between the size of GPU memory consumption and the semantic information retained in the attention map $\mathbf{A}$, we chose a $48\times 48$ size for $\mathbf{A}$ when the output resolution is $384\times 384$. We use DDIM Song et al. (2020) as our sampling method. In the experiment, $T_{1}$ in Fig.2 is set as 10, which means the SC and the STAM are employed during the first 10 timesteps; $T_{2}$ is set to 20, thus the total denoising timestep $T$ is 30. All of our experiments are conducted on a single A100 GPU.
5.2 Comparisons with Others
We evaluate the effectiveness of our method from both qualitative and quantitative perspectives. Two video diffusion baseline methods Sterling (2023); Wang et al. (2023a) are involved to confirm that our approach maintains generative performance. In addition, the SOTA pre-trained model MotionCtrl Wang et al. (2023c) is also involved in experiments to demonstrate the control ability of our proposed method.
Qualitative Results. The results of the qualitative experiments are shown in Fig.3. The base model depicted in the first two rows of the figure is ModelScope. The first row incorporates additional input boxes and utilizes the Motion-Zero framework, while in the second row, an extra prompt is added: moving from left to right. As shown in the first row left, the fish follows a trajectory that strictly adheres to the positions of the boxes. Conversely, in the second row left, the direction of the fish’s movement is not fixed, and the locations where it appears are generated randomly. In the second row right, the bear gradually rotates clockwise and fails to follow the guidance provided by the additional prompts. We can see that the model controlled by the Motion-Zero framework is capable of more effectively governing the trajectory and direction of an object’s movement which cannot be achieved by prompts alone. The base model for the third and fourth rows is ZeroScope. In the third row, both the fish and the bear move according to the conditions set by our boxes, whereas in the fourth row, they remain mostly stationary. This shows that our Motion-Zero model effectively facilitates trajectory control across various baseline models. The fifth row in the figure demonstrates the effects of the pre-trained model MotionCtrl. It can be seen that our results are comparable to those of the pre-trained model in terms of controlling the direction of the movements of the fish. In the case of generating a bear, MotionCtrl only manages to rotate the bear’s body, whereas Motion-Zero can effectively generate the bear’s movement. Note that our method additionally allows for the control of their size. Our advantage also lies in the ability to plug and play with any base model without the need for any training.
Quantitative Results with Automatic Metrics. Followed by LOVEU-TGVE competition Wu et al. (2023b), the CLIP score Hessel et al. (2021) is used to verify text-video consistency (Text Align) and inter-frame consistency (Consistency). The PickScore Kirstain et al. (2023) is used to predict user preferences of our model. We utilized five prompts, with each prompt generating ten videos to serve as test data. The results are shown in Tab.1. From the first four rows of the table, we can discern that employing our Motion-Zero framework to direct the motion trajectory of objects does not compromise the base model’s performance. Instead, the motion subject being constrained by the box results in improved semantic accuracy and temporal consistency. Upon examining the fifth row of the table, it is evident that Motion-Zero achieves an enhancement in performance over pre-trained methodology.
User Study. We recruited 5 voters from social media to evaluate 10 generated videos from the aspect of appearance, temporal consistency, and control capacity, which evaluate whether the position of the moving object is similar to the references. Voters are presented with pairs of videos in a randomized sequence without revealing the creation methods. They are then tasked with assessing each pair and indicating which video they find more favorable. The results are shown in Tab.2. The number in the table denotes the number of Motion-Zero enhanced videos preferred by the voters. v.s. ZeroScope and v.s. ModelScope are the two baseline methods supplemented with guiding prompts, for instance, moving from left to right. v.s. MotionCtrl is a comparison between ZeroScope augmented with our Motion-Zero and the standalone MotionCtrl approach with the same control trajectory. We can see that the voters prefer our methods from all aspects compared with baselines. Moreover, when compared to MotionCtrl, users demonstrated a clear preference for the control capabilities of Motion-Zero. Note that our proposed approach does not need any training. Thus, the baseline model’s inherent limitations resulted in our approach being slightly inferior to MotionCtrl in aspects of consistency.
Text Align Consistency PickScore
ZeroScope 20.05 0.89 18.15
 +Ours 22.04 0.93 20.00
ModelScope 20.30 0.91 18.56
 +Ours 24.74 0.91 19.82
MotionCtrl 20.94 0.92 19.62
Table 1: Automatic metric on baseline methods and SOTA methods. All metrics are such that higher values indicate better performance.
Appearance Consistency Control
v.s. ZeroScope 41/50 20/50 46/50
v.s. ModelScope 38/50 34/50 46/50
v.s. MotionCtrl 25/50 23/50 36/50
Table 2: User study on different motion control methods.
Figure 4: Ablation studies on different components. We use ZeroScope as our baseline model. To demonstrate the coherence of the generated object’s motion, we captured every other frame, resulting in a total of 6 frames. The prompt is A lion is walking on the field. Zoom in for the best view.
5.3 Ablation Study
We conduct a series of ablation experiments to verify the effectiveness of our modules individually.
Impact of INPM. INPM is designed to provide a prior for the initial position of moving objects, facilitating the model to more stably control the position of the object’s movement. As shown in Fig.4, the second row demonstrates the effects of removing this module. We can see that the lion is not entirely within the frame, and due to the forceful control by SC, the lion’s body undergoes significant deformation, while the lion’s head shows no movement at all. For Motion-Zero, the lion’s body is clearly generated, and the lion’s head also moves with the changes in the frame, resulting in the high overall stability of the lion. This shows that the INPM module plays a significant role in stably controlling the position of objects and generating stable videos.
Impact of SC. In this section, we validate the importance of SC. SC primarily provides auxiliary support for control over spatial positioning. As shown in the third line in Fig.4, without spatial constraints, the generated motion of the lion appears mostly outside the box. The SC is the most important module for our Motion-Zero controlling method. The reason is that SC is the only gradient-based modification to the baseline compared with other modules.
Impact of STAM. The role of the STAM is to ensure the temporal consistency of the moving object, which is to maintain the continuity of its motion as it moves. As illustrated in Fig.4, we present the effects of omitting the STAM module in the last row. It can be noted that during the movement, the lion’s legs undergo deformation and become blurred. In contrast, the first row demonstrates that the lion’s leg movements are well-generated with strong continuity, highlighting the significance of our STAM. More details are provided in the supplementary materials.
6 Discussion, Limitation and Future Works
Due to the absence of the need for additional training, our method can be applied to all pre-trained video diffusion models. However, on the other hand, its generative performance is entirely dependent on the base model. Existing video diffusion models sometimes generate relatively unstable objects, such as deformation in moving objects, holes in the background, and various issues like the inability to precisely control the attributes of generated objects with text.
Our method can currently effectively control the movement of objects in videos, but the trajectory of the movement is entirely controlled by the user, lacking semantic interaction with the video. For example, in a forest scene, if a user provides the first frame as a reference, they can use a prompt like a little rabbit goes to the stream to drink water to control a rabbit to navigate and avoid obstacles in the forest automatically. This kind of interaction between object movement and the background, as well as narrative-like semantic control, will be the direction of our future work.
7 Conclusion
In this paper, we proposed a framework Motion-Zero for controlling object motion that can be applied to any video diffusion model. Unlike previous methods that require extensive training, our method enables arbitrary motion control in a zero-shot setting. We find that the initial noise plays a significant role in object motion, leading us to introduce an initial noise prior module to achieve stable control effects. We discovered that the values in the cross-attention map significantly influence the position and appearance of the object corresponding to the prompt. Therefore, spatial constraints are designed to optimize the position of moving objects and ensure spatial consistency. Furthermore, to ensure the consistency of object motion, a shift temporal attention mechanism is developed to maintain the temporal coherence of moving objects. Extensive experiments substantiate the efficacy and generalization of our approach.